OpenStudy (anonymous):

How do I find the square root of 40 using only paper?

OpenStudy (alexandervonhumboldt2):

sqrt40=sqrt(4*10)=sqrt(2*2*10)=2sqrt10

OpenStudy (anonymous):

but how do i find the sqrt10 and so on?

OpenStudy (anonymous):

@twidget14 You could just estimate the square root of 10 based off of perfect squares. For example, 9 is a perfect square which is close to 10, and the next closest is 16. What perfect square is closest to 10? It is 9 obviously. The square root of 9 is 3, so the square root of 10 is a little over 3.

OpenStudy (radar):